Documenta Mathematica

Band 19, 2014

Yuval Z. Flicker Eisenstein Series and the Trace Formula

for GL(2) over a Function Field  1–62

Giuseppe De Nittis and Max Lein The Perturbed Maxwell Operator

as Pseudodifferential Operator  63–101

Li Chen, Jinhuan Wang Exact Criterion for Global Existence and

Blow Up to a Degenerate Keller-Segel System  103–120

Gonçalo Tabuada

E_{n}-Regularity Implies E_{n−1}-Regularity  121–139

Ivo Dell’Ambrogio, Heath Emerson, and Ralf Meyer An Equivariant Lefschetz Fixed-Point Formula

for Correspondences  141–193

Heng Xie An Application of Hermitian K-Theory:

Sums-of-Squares Formulas  195–208

Timo Richarz A New Approach

to the Geometric Satake Equivalence  209–246

Manabu Murata On Quadratic Diophantine Equations in Four

Variables and Orders Associated with Lattices  247–284

Wilhelm Plesken, Thomas Bächler Counting Polynomials for Linear Codes,

Hyperplane Arrangements, and Matroids  285–312

Mikhail Ershov, Wolfgang Lück The First L^{2}-Betti Number

and Approximation in Arbitrary Characteristic  313–331

Nicola Ciccoli, Fabio Gavarini A Global Quantum Duality Principle

for Subgroups and Homogeneous Spaces  333–380

Christian Wuthrich On the Integrality of Modular Symbols and

Kato’s Euler System for Elliptic Curves  381–402

iv